OpenStudy (anonymous):

To graph the equation y= -1/2x - 3 using the slope and y-intercept, you would have to do which of the following?

OpenStudy (senkai):

Is there a multiple choice to choose from?

OpenStudy (anonymous):

Graph the y-intercept, 3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, 3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points.

OpenStudy (anonymous):

@Senkai

OpenStudy (anonymous):

from the y intercept, which is -3, you come down 1 and go to the right 2, then down 1 and to the right 2. Because the slope (rise/run) is -1/2.

OpenStudy (anonymous):

yep so awnser choice C?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

i got anthor one @kelliegirl33

OpenStudy (anonymous):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes

OpenStudy (anonymous):

what are the points ?

OpenStudy (anonymous):

set #1 (4,4) and (2,6) and set # 2 (2,4) and (4,2) @kelliegirl33

OpenStudy (anonymous):

The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

are you familiar with the slope formula ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so lets find the slopes.... set 1 : (4,4) x1 = 4, and y1 = 4 (2,6) x2 = 2 and y2 = 6 slope(m) = (y2 - y1)/(x2 - x1) m = (6 - 4) / (2 - 4) m = 2/-2 m = -1 (slope of set 1) set 2 : (2,4) x1 = 2 and y1 = 4 (4,2) x2 = 4 and y2 = 2 m = (2 - 4) / (4 - 2) m = -2/2 m = -1 (slope of set 2) so what do you think the answer is ?

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

the awnser is C?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

possibly....whatcha got

OpenStudy (anonymous):

Using complete sentences, explain how to graph y = negative one fourthx + 8 using the slope-intercept method.

OpenStudy (anonymous):

y = -1/4x + 8 In this form, y = mx + b, the slope is m and the y intercept is b. Therefore, the slope is -1/4 and the y intercept is 8 Start at the y intercept and plot your point on 8. From that point, come down 1 and to the right 4, then down 1 and to the right 4. Because the slope is -1/4 (rise/run), since the 1 is negative, you come down 1, and since the 4 is positive, you go to the right 4.

OpenStudy (anonymous):

omg thanks so muchh!!!!:)

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

cool....I was glad I could help